If heidi's current age in years is multiplied by five and two years is subtracted from the answer the result is equal to her dad's age. In 8 years time the ages will be as such that heidi's dad will be three times as old as heidi. How old is heidi now?
i tried
x X 5 - 2 = Dads age
then i was thinking
x X 5 - 2 = 3(x X5 -2 )
I am stuck

Answer by Paul(988)

5x-2=y --->y=5x-2
In eight years her dad's age = y+x
In eight years her age = x+8

3(x+8)=y+8
SUbsitute:
3x+24=5x-2+8
3x+24=5x+6
-2x=-18
x=9

y=5(9)-2
y=43

Hence, Heidi is 9 years old and her dad is 43 years old.
Paul.